OpenStudy (skyleekinton):

3/4x−1/2(2−1/2x)=1/3 What is the value of x? Enter your answer in the box. x= @leenathan @Jack_Prism @Jason4Jesus

OpenStudy (leenathan):

0.59628479

OpenStudy (skyleekinton):

Is that the real answer lee

OpenStudy (leenathan):

used this https://www.mathway.com/Algebra

OpenStudy (welshfella):

Please don't give direct answers!

OpenStudy (leenathan):

oh ok sorry

alones (alones):

Otherwise explain it Lee

OpenStudy (leenathan):

http://prnt.sc/dcu8jf

alones (alones):

Or simply simplify \[ \frac{ 3 }{ 4 }-\frac{ 1 }{ 2 }(2-\frac{ 1 }{ 2 }x=\frac{ 1 }{ 3}\] to this \[x-1=\frac{ 1 }{ 3 }\] Than add \(1\) to each side.

alones (alones):

\[x−1+1=\frac{ 1 }{ 3 }+1\] For which I got \(x=\frac{ 4 }{ 3 }\)

OpenStudy (skyleekinton):

Thx Alone

OpenStudy (welshfella):

the question is not clear The way its written can be interpreted in different ways